# Homological shifts of a complementary edge ideal

### v1

@@ 4.3. Cycles
For $n=2m+1$ and $i=2m$, substituting these values gives $\left\lceil\frac{n}{2}\right\rceil=m+1$, so the left bound simplifies to $\max\{2m-(m+1)+1,\ 0\}=\max\{m,\ 0\}=m$. The right bound is $\left\lfloor\frac{2m}{2}\right\rfloor=m$, hence $k=m$ is the only possible value. Thus, the result follows from Formula (5).
The proof of (2) is analogous.    ∎
@@ Acknowledgments
The first author is supported by the National Natural Science Foundation of China (No. 11971338). The third author is supported by the Natural Science Foundation of Jiangsu Province (No. BK20221353) and the National Natural Science Foundation of China (No. 12471246).
Data availability statement
The data used to support the findings of this study are included within the article.
Conflict of interest
The authors declare that they have no competing interests.

### v5

@@ 5.3. Cycles
For the case where $n=2m+1$ and $i=2m$, we have $\left\lceil\frac{n}{2}\right\rceil=m+1$. The left-hand bound of the above inequality simplifies to
$\max\{2m-(m+1)+1,\ 0\}=\max\{m,\ 0\}=m,$
while the right-hand bound is $\left\lfloor\frac{2m}{2}\right\rfloor=m$. Therefore, $k=m$ is the only feasible value of $k$, so i2k=0i-2k=0. The equality $\mathrm{HS}_{2m}(I^{s})=\alpha^{m}\cdot I^{s-m}$ thus follows immediately from Formula (6), and it then follows that $\mathrm{HS}_{2m}(\mathcal{R}(I))$ is generated precisely in degree $m$.
The proof of part (2) is analogous to that of part (1).
@@ Acknowledgments
The first author is supported by the National Natural Science Foundation of China (No. 11971338). The third author is supported by the Natural Science Foundation of Jiangsu Province (No. BK20221353) and the National Natural Science Foundation of China (No. 12471246). The authors wish to express their gratitude to the computer algebra system [8], which provided a wealth of examples that facilitated the development of our research ideas and the verification of our results.